Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 582     Entry time: 28 Nov 2018, 10:18
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Status report after dispatch meeting 09.30 

SPB

  •  commissioning ongoing
  • Radiation test was not successful since the beam diameter was smaller than expected. 

FXE

SQS

SCS

MID

  • Loop 7 will run by today evening
  • Mono commissioning issue: the pico motors position is not reproducible. ==> the issue solved by itself. Tests ongoing.
  • AGIPD commissioning is ongoing. 

HED

XRO

XPD

VAC

AE

  • In sase3 the y axis of the nano motion system (sample environment) does not work properly uinder vacuum. JTolkien and C Dieter are on the issue.

EET

ITDM

 

CAS

  • Release tests ongoing. Release candidate being tested with AE.
  • OCD for SASE3 users experiment has been set up

AoB

ELOG V3.1.4-7c3fd00